OpenStudy (anonymous):

a speedboat traveled 2 hours with a 4km/h current before turning around. the return trip against the same current took 3 hours. find the speed of the boat in still water

OpenStudy (amistre64):

s = d/t is usually used to aid in this word problem

OpenStudy (anonymous):

ok im just trying to figure if i got it right after solving it wat did u get

OpenStudy (amistre64):

total speed = boatspeed + currentspeed ; STILL WORKING ON IT LOL

OpenStudy (anonymous):

2(x+4)=3(x-4)

OpenStudy (amistre64):

they really should move the caps lock key to a new place

OpenStudy (anonymous):

its cool take ur time u r the one helping me lol i kno right

OpenStudy (amistre64):

imran; has a good setup

OpenStudy (anonymous):

speed in still water is 20

OpenStudy (anonymous):

yeah its just i never get past that part i mean i kno it easy but........

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

welcome

OpenStudy (amistre64):

think of it as: distance = distance time(speed) = time(speed)

OpenStudy (anonymous):

well put amistre

OpenStudy (anonymous):

ok easier ty

OpenStudy (amistre64):

2(boat speed + current) = 3(boat speed - current)